OpenStudy (anonymous):

Show your work: (x-2)/4 + x/3 = 3

OpenStudy (ash2326):

@Elysse2012 did you try this?

OpenStudy (anonymous):

yeah.... i did this (12) ((x-2)/4) + (12)(x/3) = (12)(3) because 12 is the LCM and i was wondering if what i got was right. I got 12x-6/3 + 12/4 = 36 which is 4x-2+3 = 36 then 4x+1 = 36..... but that's not right i dont think.

OpenStudy (ash2326):

LCM is 12 that's correct now we have \[\frac{x-2}{4}+\frac{x}3=3\] multiply by 12 \[12\times \frac{x-2}{4}+12\times\frac x3=12\times 3\] \[\cancel{12}3\times \frac{x-2}{\cancel 4}+\cancel{12}4\times\frac x {\cancel 3}=12\times 3\] we get \[3(x-2)+4x=36\] \[3x-6+4x=36\] Can you solve now?

OpenStudy (anonymous):

whered the 12 3 and the 12 4 come from?

OpenStudy (ash2326):

12 got cancelled by 4 and then we get 3

OpenStudy (ash2326):

@Elysse2012 ?

OpenStudy (anonymous):

ok. i get what youre saying now. sigh.... I hope this all clicks soon :/

OpenStudy (ash2326):

it'll soon. Did you understand this?

OpenStudy (anonymous):

yeah. thanks!